OpenStudy (anonymous):

a rectangular box with a square base and open top is to be made from 1200cm^2 of material. 1. Find the dimensions of the largest box that can be made. 2.Use the second derivative test to show that the dimensions in 1 give a maximum volume.

OpenStudy (anonymous):

So, The Total Surface of the box is 1200cm^2

OpenStudy (anonymous):

the total of the box surface area is \[x ^{2}+4xh\] so do i equate \[1200=x ^{2}+4xh\]

OpenStudy (anonymous):

Let y^2 be the Area of the square base and h be the height of the box then, y^2+4y*h=1200

OpenStudy (anonymous):

h=(1200-y^2)/4y......i

OpenStudy (anonymous):

volume is V=hbl then do i use the above equation to find h?

OpenStudy (anonymous):

we seem to be thinking the same thing at the same time.

OpenStudy (anonymous):

Now, VOLUME = AREA OF BASE * height

OpenStudy (anonymous):

V=y^2 *h V=y^2 *(1200-y^2)/4y

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

V=y^2 *(1200-y^2)/4y V=300y - 1/4 y^3

OpenStudy (anonymous):

Now, Find the maximum value of V

OpenStudy (anonymous):

how did you get V=300y - 1/4 y^3? i get x(1200-x^2)/4

OpenStudy (anonymous):

That is same thing.

OpenStudy (anonymous):

wait i see now my bad

OpenStudy (anonymous):

only we are using different variable.

OpenStudy (anonymous):

Find V'

OpenStudy (anonymous):

so i find V'(x)=0 for maximum?

OpenStudy (anonymous):

so \[300-\frac{ 3y ^{2} }{ 4 }=0\]

OpenStudy (anonymous):

at x = 20 we will have a maximum?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Now find height when x=20

OpenStudy (anonymous):

so volume will be 4000cm^3

OpenStudy (anonymous):

YEP

OpenStudy (anonymous):

does y aply for H, B and L?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

nvm. I have a height of 10cm and a width and breadth of 20cm?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

and the second part?

OpenStudy (anonymous):

FIND v''

OpenStudy (anonymous):

-3/2x

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Now, when x=20 V''=-3/2 *20 =-30

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

Which is negative, Thus V is maximum when x=20

OpenStudy (anonymous):

so if the V'' is negative its maximum at point x and if it is positive then it is a minimum at point x?

OpenStudy (anonymous):

is that just in general or is it for a fixed number?

OpenStudy (anonymous):

U find x-coordinate of the stationary point FROM V'=0

OpenStudy (anonymous):

Now, if for that x-coordinate , V'' is negative then the stationary point is maximum is positive then minimum and is 0 then neither maximum nor minimum.

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

WElcome